Sheffield Corporation entered into a lease agreement on January 1, 2017, to provide Pharoah Company with a piece of machinery. The terms of the lease agreement were as follows.
1. The lease is to be for 3 years with rental payments of $13,700 to be made at the beginning of each year.
2. The machinery has a fair value of $65,000, a book value (depreciable base for the lessor) of $40,000, and an economic life of 8 years.
3. At the end of the lease term, both parties expect the machinery to have a residual value of $25,000, none of which is guaranteed.
4. The lease does not transfer ownership at the end of the lease term, does not have a bargain purchase option, and the asset is not of a specialized nature.
5. The implicit rate is 4%, which is known by Dawkins.
6. Collectibility of the payments is probable.


Click here to view the factor table.

Collapse question part

(b)

Prepare the amortization schedules Pharoah will use over the lease term.
PHAROAH COMPANY
Lease Amortization Schedule
Annuity-Due Basis
Date Annual Payment Interest
on Liability
Reduction of
Lease Liability
Lease Liability
1/1/17 $

1/1/17 $

$

$

1/1/18

1/1/19

Click if you would like to Show Work for this question:

Open Show Work
Answer #1
Lease Amortisation Schedule
Date Beg Bal of Lease Liability Interest on Lease Liability Reduction of Lease
Liability
Annual Lease Payment End Bal of Lease Liability
a b=a*4% c=d-b d a-c
1/1/2017 $39,540 $39,540
1/1/2017 $39,540 $13,700 $25,840
1/1/2018 $25,840 $1,033.60 $12,666 $13,700 $13,174
1/1/2019 $13,174 $526.94 $13,173 $13,700 $1
Computation of Present Value of Lease Rental
Annual Lease Rental $13,700
Cumm PVAF @ 4% for 0-2 2.8861
PV of Lease Rental
(13700X2.8861)
$39,540